# Fading into darkness: A weak mass ejection and low-efficiency fallback accompanying black hole formation in M31-2014-DS1

### v1

@@ IV.1 A low kinetic energy mass ejection
The progenitor of M31-2014-DS1 was classified as a warm YSG based on multi-wavelength SED modeling from m (De et al., 2026) – indicating a photospheric radius of R⊙ and a surrounding dust shell with an inner radius of au. In contrast, the remnant is enshrouded in a dust shell formed at much closer radii, confirming new dust formation from material ejected during the disappearance, as suggested by De et al. (2026). However, unlike De et al. (2026) who assumed that the majority of the ejected gas had condensed into dust with dust-to-gas mass ratio as in the interstellar medium (), our measurements suggest . This indicates that the ejected material has not completely condensed into dust – explaining the increasing dust optical depth between as dust formation continues.
@@
We estimate the required kinetic energy () of the explosion to explain the presence of M⊙ of molecular gas at au333Given that the molecular gas likely associated with ejected material is confined to au, it is possible that the majority of the outer dust is leftover from the progenitor dust shell.. We first highlight that the velocity of the molecular slab ( km s-1) with the inferred gas mass indicates erg as an approximate order of magnitude estimate. To further establish the estimate with quantitative modeling, we take the YSG progenitor model of De et al. (2026)444The models are available at https://github.com/dekishalay/M31-2014-DS1/ and at Zenodo (De, 2025). and inject low energy shocks into the stellar envelope. As in De et al. (2026), we adopt a simple, physically motivated model in which the ejecta are characterized by a power-law radial velocity profile defined as,
@@
where is the radius within the stellar envelope, parameterizes the radial distribution of kinetic energy, and is a normalization that is set by the total input kinetic energy (, which is in units of erg). The total mass of the outer H-rich envelope is M⊙ , which has a gravitational binding energy of erg.Fa
@@
Appendix A describes our numerical method for computing the radial evolution of the envelope mass – demonstrating the competing effects of mass ejection and fallback at different values of . We use this method to compute the radial density profile of material as a function of time – which is sensitive to as well as the assumed . Although the exact time of collapse is not well constrained for M31-2014-DS1 due to the lack of a detected optical outburst (De et al., 2026), the MIR brightening and optical disappearance of the source suggest that the JWST observations were obtained at a phase of d after collapse. We show the calculated cumulative mass profile of the ejected mass and fallback material555We use d as the estimated phase of the observation, noting that the radial density profile moves laterally in the figure by % within the range of possible times since the ejected material is traveling freely at this phase. for different values of and in Figure 4. A higher kinetic energy unbinds a larger fraction of the envelope and places the ejected gas at a larger radii at a fixed time after explosion. A higher leads to a larger amount of energy in the high velocity layers such that a larger fraction of the envelope falls back at a fixed energy.
@@
Figure 4 shows that the inferred location of the gas layer is consistent with ejection energies of erg666The low energy is consistent with the complete lack of an optical outburst prior to the fading (De et al., 2026; Antoni et al., 2025).. While an ejection energy that is lower produces too much fallback (leaving little circumstellar material at au), a higher value ejects the material at too high velocities that reaches larger radii ( au) in this time span (see also Appendix A). We highlight that the physical cause of the energy injection is not directly constrained by these observations; in the failed SN scenario, we expect this to be a combination of the neutrino mass-loss driven shock (Nadezhin, 1980; Lovegrove and Woosley, 2013) and energy injection due to the angular momentum barrier encountered by infalling material (the majority of which falls back at early times; see Figure 6) from the outer envelope (Quataert et al., 2019; Antoni and Quataert, 2023; Antoni et al., 2025).
@@ IV.2 Accretion-powered heating from inefficient fallback
While impulsive energy injection into the stellar envelope is agnostic to the physical scenario driving it, the strongest evidence for the failed SN hypothesis for M31-2014-DS1 (as well as NGC 6946-BH1) lies in its continued bolometric fading to substantially fainter levels than the progenitor star. As argued in De et al. (2026); Kochanek (2024), the bolometric fading of observed until 2022 suggests a terminal collapse of the stellar core, implying the birth of a BH from the core-collapse. While the analysis of De et al. (2026) only include MIR data to m, the new JWST observations rule out the presence of a cold dust component at m (which would manifest a long wavelength excess at m, unlike the observed fading SED at m in Figure 1) that could hide the stellar luminosity due to dust obscuration. Figure 4 shows the bolometric light curve combining published measurements with the latest JWST observations – further confirming the continued bolometric fading of the source and the terminal disappearance of the star.
@@
In the failed SN scenario, the late-time luminosity is powered by accretion of marginally bound material that is gradually accreted into the nascent BH – noting that the majority of the material (%) is expected to eventually be expelled due to the random angular momentum barrier near the BH horizon (Quataert et al., 2019; Antoni and Quataert, 2022, 2023). In the case of purely spherical infall, the emergent luminosity would be much lower (Faran and Quataert, 2025). De et al. (2026) argued that the bolometric fading of both M31-2014-DS1 and NGC 6946-BH1 requires that the progenitor be depleted of H in its envelope to explain the fading over timescales of yr. This is because the sustained ( d), highly super-Eddington accretion rates expected for red supergiant progenitors with massive ( M⊙ ) H-envelopes are expected to produce a near-constant luminosity capped at the Eddington luminosity ( erg s-1 for a M⊙ BH) for centuries – followed by fading on similar timescales (Fernández et al., 2018; Coughlin et al., 2018; Faran and Quataert, 2025). The depleted ( M⊙ ) H-envelope was independently confirmed by the secure YSG progenitor classification for M31-2014-DS1 and supported by data for NGC 6946-BH1 (Humphreys, 2019).
Adopting the constraints on the ejection energy ( erg) from the gas kinematics, we find that the fallback model presented in De et al. (2026) leads to much brighter emission and slower fading at a phase of yr than observed (Figure 4). However, the model of De et al. (2026) is limited in that i) it assumed a constant, nominal radiative efficiency of and ii) it estimated the fraction of material () that is truly accreted based on an approximate analytic prescription based on the relative (random) angular momentum of infalling material to that at the BH horizon.
@@
In reality, additional effects such as mechanical and radiative feedback from the BH accretion can further reduce the fraction of material that is accreted (Antoni and Quataert, 2023; Antoni et al., 2025). As shown in Figure 4, the observed bolometric fading is consistent with explosion energies of erg if only a small fraction of the mass is accreted compared to that indicated by the angular-momentum based analytic prescription. We therefore find that either the combination of the angular momentum barrier and mechanical/radiative feedback on the infalling gas leads to highly inefficient accretion of the fallback material such that only % of the total envelope mass is accreted in the long term or that the radiative efficiency is % – much lower than accreting BH X-ray binary systems (McClintock and Remillard, 2006).
@@
Despite the high bolometric luminosity emerging from the accreting BH ( erg s-1 emerging entirely in the IR band), no X-ray emission (expected to emerge from close to the BH horizon) was detected to erg s-1. We attribute this to extremely high inferred column density of H-gas from the ejected material – which is directly measured as and alone sufficient to obscure the X-ray emission by a reprocessing layer (De et al. 2026; see their Figure S10) – in addition to the fallback material at R⊙ . While associated with the ejecta should decline with time as , De et al. (2026) approximated that the fallback component becomes dominant () at late-times. Given the low inferred ejection energy, the nominal model for the fallback column density (based on the mass density near the turn-around radius; De et al. 2026) would indicate an even higher contribution of the fallback material to at this phase ().
@@ V Summary
The remnant is enshrouded in an optically thick dust shell with , and has become progressively redder since the 2022 observations reported in De et al. (2026). Comparing to the progenitor properties, we show formation of M⊙ of new dust at a radius of au.
@@
The source continues to fade dramatically in bolometric light, reaching % of the progenitor luminosity in 2024. We use the continued bolometric fading as evidence for a terminal mass ejection that led to the formation of a BH and a slowly fading remnant powered by fallback accretion.
@@
Appendix A Ejection and fallback of material

### v2

@@ IV.1 A low kinetic energy mass ejection
The progenitor of M31-2014-DS1 was classified as a warm YSG based on multi-wavelength SED modeling from m (D26) – indicating a photospheric radius of R⊙ and a surrounding dust shell with an inner radius of au. In contrast, the remnant is enshrouded in a dust shell formed at much closer radii, confirming new dust formation from material ejected during the disappearance, as suggested by D26. However, unlike D26 who assumed that the majority of the ejected gas had condensed into dust with dust-to-gas mass ratio as in the interstellar medium (), our measurements suggest . This indicates that the ejected material has not completely condensed into dust – explaining i) the increasing dust optical depth between as dust formation continues and ii) the differing temperatures of the gas and dust species (Table 1) as the material has not reached equilibrium. However, we caution that is subject to uncertainties in and precise dust characteristics (degeneracies in grain size and composition).
@@
We estimate the required kinetic energy () of the explosion to explain the presence of M⊙ of molecular gas at au555Given that the molecular gas likely associated with ejected material is confined to au, it is possible that the majority of the outer dust is leftover from the progenitor dust shell.. We first highlight that the velocity of the molecular slab ( km s-1) with the inferred gas mass indicates erg as an approximate order of magnitude estimate. To further establish the estimate with quantitative modeling, we take the YSG progenitor model of D26666The models are available at https://github.com/dekishalay/M31-2014-DS1/ and at Zenodo (De, 2025). and inject low energy shocks into the stellar envelope. As in D26, we adopt a simple, physically motivated model in which the ejecta are characterized by a power-law radial velocity profile defined as,
@@
where is the radius within the stellar envelope, parameterizes the radial distribution of kinetic energy, and is a normalization that is set by the total input kinetic energy (, which is in units of erg). The total mass of the outer H-rich envelope is M⊙ , which has a gravitational binding energy of erg.
@@
Appendix B describes our numerical method for computing the radial evolution of the envelope mass – demonstrating the competing effects of mass ejection and fallback at different values of . We use this method to compute the radial density profile of material as a function of time – which is sensitive to as well as the assumed . Although the exact time of collapse is not well constrained for M31-2014-DS1 due to the lack of a detected optical outburst (D26), the MIR brightening and optical disappearance of the source suggest that the JWST observations were obtained at a phase of d after collapse. We show the calculated cumulative mass profile of the ejected mass and fallback material777We use d as the estimated phase of the observation, noting that the radial density profile moves laterally in the figure by % within the range of possible times since the ejected material is traveling freely at this phase. for different values of and in Figure 4. A higher kinetic energy unbinds a larger fraction of the envelope and places the ejected gas at a larger radii at a fixed time after explosion. A higher leads to a larger amount of energy in the high velocity layers such that a larger fraction of the envelope falls back at a fixed energy.
@@
Figure 4 shows that the inferred location of the gas layer is consistent with ejection energies of erg888The low energy is consistent with the complete lack of an optical outburst prior to the fading (D26, Antoni et al. 2025).. While an ejection energy that is lower produces too much fallback (leaving little circumstellar material at au), a higher value ejects the material at too high velocities that reaches larger radii ( au) in this time span (see also Appendix B). We highlight that the physical cause of the energy injection is not directly constrained by these observations; in the failed SN scenario, we expect this to be a combination of the neutrino mass-loss driven shock (Nadezhin, 1980; Lovegrove and Woosley, 2013) and energy injection due to the angular momentum barrier encountered by infalling material (the majority of which falls back at early times; see Figure 7) from the outer envelope (Quataert et al., 2019; Antoni and Quataert, 2023; Antoni et al., 2025).
@@ IV.2 Accretion-powered heating from inefficient fallback
While impulsive energy injection into the stellar envelope is agnostic to the physical scenario driving it, the strongest evidence for the failed SN hypothesis for M31-2014-DS1 (as well as NGC 6946-BH1) lies in its continued bolometric fading to substantially fainter levels than the progenitor star. As argued in D26 and Kochanek (2024), the bolometric fading of observed until 2022 suggests a terminal collapse of the stellar core, implying the birth of a BH from the core-collapse. While the analysis of D26 only include MIR data to m, the new JWST observations rule out the presence of a cold dust component at m that could hide the stellar luminosity due to dust obscuration. Figure 4 shows the bolometric light curve combining published measurements with the latest JWST observations – further confirming the continued bolometric fading of the source and the terminal disappearance of the star.
@@
In the failed SN scenario, the late-time luminosity is powered by accretion of marginally bound material that is gradually accreted into the nascent BH. However, the majority of the material (%) is expected to be eventually expelled due to the random angular momentum barrier near the BH horizon (Quataert et al., 2019; Antoni and Quataert, 2022, 2023). In the case of purely spherical infall, the emergent luminosity would be much lower (Faran and Quataert, 2025). D26 argued that the bolometric fading of both M31-2014-DS1 and NGC 6946-BH1 requires that the progenitor be depleted of H in its envelope to explain the fading over timescales of yr. This is because the sustained ( d), highly super-Eddington accretion rates expected for RSG progenitors with massive ( M⊙ ) H-envelopes are expected to produce a near-constant luminosity capped near the Eddington luminosity ( erg s-1 for a M⊙ BH) for centuries. As the mass accretion rate drops below the Eddington rate after this time, the luminosity would fade with time as on similar timescales (over many centuries for RSG progenitors; Fernández et al. 2018; Coughlin et al. 2018; Faran and Quataert 2025). The depleted ( M⊙ ) H-envelope was independently confirmed by the secure YSG progenitor classification for M31-2014-DS1 and independently supported by data for NGC 6946-BH1 (Humphreys, 2019).
@@
Adopting the constraints on the ejection energy ( erg) from the gas kinematics, we find that the fallback model presented in D26 leads to much brighter emission and slower fading at a phase of yr than observed (Figure 4; again decaying with time as at these late phases). However, the model of D26 assumed i) a constant, nominal radiative efficiency of and ii) an estimate of the fraction of material () that is truly accreted based on an approximate analytic prescription based on the relative (random) angular momentum of infalling material to that at the BH horizon. In reality, additional effects such as mechanical and radiative feedback from the BH accretion can further reduce the fraction of material that is accreted (Antoni and Quataert, 2023; Antoni et al., 2025). As shown in Figure 4, the observed bolometric fading is consistent with explosion energies of erg if only a small fraction of the mass is accreted compared to that indicated by the angular-momentum based analytic prescription. We therefore find that either the combination of the angular momentum barrier and mechanical/radiative feedback on the infalling gas leads to highly inefficient accretion of the fallback material such that only % of the total envelope mass is accreted in the long term or that the radiative efficiency is % – much lower than accreting BH X-ray binary systems (McClintock and Remillard, 2006).
@@
Despite the high bolometric luminosity emerging from the accreting BH ( erg s-1 emerging entirely in the IR band), no X-ray emission (expected to emerge from close to the BH horizon) was detected to erg s-1. We attribute this to extremely high inferred column density of H-gas from the ejected material – which is directly measured as and alone sufficient to obscure the X-ray emission by a reprocessing layer (D26; see their Figure S10) – in addition to the fallback material at R⊙ . While associated with the ejecta should decline with time as , D26 approximated that the fallback component becomes dominant () at late-times. Given the low inferred ejection energy, the nominal model for the fallback column density (based on the mass density near the turn-around radius; D26) would indicate an even higher contribution of the fallback material to at this phase ().
@@ V Summary
The remnant is enshrouded in an optically thick dust shell with , and has become progressively redder since the 2022 observations reported in D26 Comparing to the progenitor properties, we show formation of M⊙ of new dust at a radius of au.
@@
The source continues to fade dramatically in bolometric light, reaching % of the progenitor luminosity in 2024. Together with the yr archival history of constant brightness, we use the continued bolometric fading as evidence for a terminal mass ejection that formed a BH and a slowly fading remnant powered by fallback accretion.
@@
Appendix A Parameter posterior distributions from DUSTY
We fit the MIR SED of M31-2014-DS1 using the DUSTY code, as described in Section III. Here, we show the posterior distributions of the resulting fit in Figure 6. The parameters do not show significant degeneracies in the distribution. While some parameters (e.g. the total flux ) are very well constrained, the grain size is not; however, it does not affect our analysis.
Appendix B Ejection and fallback of material